OpenStudy (shatha):

hi .. i am a new here and i need your help in this question.. limx__>0 √(x+25)-5/1-√x+1

OpenStudy (anonymous):

x__>-sqrt(x)-4

OpenStudy (kira_yamato):

Can you use the equation function, please.... I appreciate it...

OpenStudy (anonymous):

(x__) > -4 - Sqrt[x

OpenStudy (shatha):

are u sure??

OpenStudy (jamesj):

Ok, you want \[\lim_{x \rightarrow 0} \frac{ \sqrt{x+25} - 5}{1 - \sqrt{x+1}}\] yes?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

yes

OpenStudy (jamesj):

multiply top and bottom by \[\sqrt{x+25} + 5\] and rationalize the numerator

OpenStudy (jamesj):

then multiply top and bottom by \[1 + \sqrt{x+1}\] and rationalize

OpenStudy (jamesj):

you should then be able to cancel out an x from top and bottom and then be able to take the limit

OpenStudy (shatha):

thanx alot ..OMG i think i did it wrong in the quiz

OpenStudy (shatha):

=(